# Exploring the Impact of HAPS-RIS on UAV-Based Networks: a Novel Network Architecture

### v2

@@ V Conclusion
This paper introduces a network architecture that leverages HAPS-RIS to assist UAV-based wireless networks. A significant challenge in these networks is the limited number of UAVs, which restricts full coverage and leaves some users unsupported. To address this, we propose a hierarchical bilevel optimization framework consisting of a leader and follower problem. The leader problem aims to obtain the zone boundary and a practical RIS phase shift design that maximizes the number of users supported by the HAPS-RIS while ensuring that their rate requirements are met. This is achieved through our proposed relaxation method and the algorithm based on the RIS clustering concept. The follower problem minimizes the number of UAVs needed while still meeting the rate requirements of the users in the UAV zone.

### v3

@@ V Conclusion
This paper proposes a HAPS-RIS-assisted network architecture to enhance UAV-based wireless coverage, addressing the challenge of limited UAV availability. A hierarchical bilevel optimization framework is designed, where the leader problem determines the zone boundary and practical RIS phase shifts to maximize HAPS-RIS-supported users under rate constraints. This is achieved using a relaxed phase design and the proposed DyRaZARC technique. The follower problem minimizes the number of UAVs required to cover remaining users by applying the proposed geometry-informed machine learning KADUS technique, which adapts $k$-means clustering to communication constraints and geometric initialization. Our study reveals that increasing the number of RIS elements significantly reduces the number of required UAVs, enables full user coverage even at high data rates, and supports scalable NTN deployment. This study opens several avenues for future research. One direction is to extend the model to mobile UAV trajectories with energy and path planning constraints. Another is to integrate sensing capabilities into the proposed network architecture, enabling intelligent, context-aware NTN deployment through integrated sensing and communication (ISAC).